OpenStudy (anonymous):

The height of a triangle is 4 m less than its base. The area of the triangle is 48 m2. Find the length of the base. A. 7 m B. 8 m C. 11 m D. 12 m

OpenStudy (nurali):

The area of a triangle equals 1/2 times base times height Let x represent the triangle's base 48 = 1/2 * x * (x-4) 96 = x^2 -4x x^2 -4x -96 = 0 factor the equation (x-12) * (x+8) = 0 x=12, x=-8 correct solution is the base equals 12 meters

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

A B C or D?

OpenStudy (anonymous):

He just said 12m, try reading the explanation instead of looking for just the answer...

OpenStudy (anonymous):

ok thenks sorry i was just skimming i'm kind of in a hurry

OpenStudy (nurali):

My Pleasure.